OpenStudy (anonymous):

6(x + 3)2 = 15(x + 3) - 6 Which of the following is a solution for x?

OpenStudy (anonymous):

Is it \[ 6(x+3)^2=15(x+3)-6\]

OpenStudy (anonymous):

yes thats what i put @BlackLabel

OpenStudy (the_fizicx99):

So its being raised to a power? @kaylahhh3

OpenStudy (anonymous):

@tHe_FiZiCx99 yes I guess, my options are x = -1 x = -5 x = 5/2 x = -3

OpenStudy (shamil98):

\[\huge 6(x+3)^2 = 15(x+3) - 6\] \[\huge 6x^2 + 36x + 54 = 15x +39\] \[\huge 6x^2 + 21x + 15\] Using the quadratic formula.. \[\huge x = \frac{ -21 \pm \sqrt{21^2 - 4(6)(15)} }{ 2(6) }\] \[\Huge x = \frac{ -21 \pm \sqrt {441-360} }{ 12 }\]

OpenStudy (shamil98):

\[\Huge x = \frac{ -21 \pm \sqrt{81} }{ 12 }\] \[\Huge x = \frac{ -21 \pm 9 }{ 2 }\]

OpenStudy (shamil98):

\[\Huge x = \frac{ -21 + 9 }{ 12 } = -1\] \[\Huge x = \frac{ -21 - 9 }{ 12 } = \frac{ -30 }{ 12 } = \frac{ -5 }{ 2 }\] \[\Huge So~x= -1, x= -\frac{ 5 }{ 2 }\]

OpenStudy (shamil98):

That was a typo in the second part \[\huge x = \frac{ -21 \pm 9 }{ 12 }\] not over 2 *